The energy of an electron in the first Bohr orbit of H-atom is -13.6 eV. The possible energy value(s) of the excited state(s) for electrons in Bohr orbits of hydrogen is (are)
a. -3.4 e V
b. -4.2 e V
c. -6.8 e V
d. +6.8e V

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. Energy of electron in H-atom is determined by the expression E_n = 13.6/n^2 eV where, n = 1, 2, 3,…. In excited states, E_2 = - ( 13.6)/4 = - 3.4 eV E_3 = ( 13.6)/9 = - 1.51 eV etc.